OpenStudy (anonymous):

What is the percent by mass of NaHCO3 in a solution of 40.0g NaHCO3 dissolved in 800.0mL of H2O?

OpenStudy (aaronq):

whats the total mass of the solution?

OpenStudy (anonymous):

i think 1st of all find the mass of solvent that is H2O with the help of density, density of water=1 g/ml mass of H2O= d X v = 1 X 800= 800 g of H2O now, total mass of solution = mass of solvent + mass of solute= 800 + 40 = 840 %age of NaHCO3=40/840 X 100 = 4.76 % by mass